What is the value of x will make de//ab,in the given fig . Ad=8x+9,cd=x+3,be=3x+4,ce=x

Grade:10

1 Answers

Arun
6 years ago
By thales theorem
We can write
CD/CA = CE/CB
(x+3)/(9x+12) = x/(4x+4)
On cross multiply
4x² + 16x +12 = 9x²+12 x
5x²-4x -12 = 0
5x²-10x +6x -12 = 0
x = -6/5 and 2